MA 08 A 17WNCMP CHE-121-20102 Week 4 Week 4Qulz What mass of HCI (30.4.0 gmo) is present in 36.2 Lof a o 0827 Msolution of HCI? red Select one: 0.00305 g 15933 g 0.0822g 4.29 x 10 g 109 g
Answer #1

Molarity is defined as moles / litre of solution. Moles of HCl / Litre of solution = 0.0827. We have 36.2 L solution. Therefore moles of HCl in the solution = 0.0827 * 36.2 = 3 moles.

Therefore mass of HCl = 3 * 36.4 = 109.2 g